OpenStudy (anonymous):

Medal and Fan! <3 Which of the following is a solution of x2 - 7x = -5? A) 7 plus or minus the square root of 69 divided by two B) 7 plus or minus the square root of 29 divided by two. C) negative 7 plus or minus the square root of 29 divided by two D) negative 7 plus or minus the square root of 69 divided by two.

OpenStudy (xmissalycatx):

Do you know how to solve the equation first of all?

OpenStudy (xmissalycatx):

By using the quadratic formula..

OpenStudy (anonymous):

I tried that. x^2-7x=-5 so we would first subtract 7x by both sides making it x^2=7x-5 then we would divide by x^2 but im just not getting where we are getting the square roots in all of this. @xMissAlyCatx

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

no okay so it would be x=-b sqrt of b^2-4ac / 2a

OpenStudy (xmissalycatx):

That's the quadratic formula, yes.

OpenStudy (anonymous):

okay so everything would equal to 0

OpenStudy (anonymous):

x^2-7x+5=0?

OpenStudy (xmissalycatx):

I really hate to tell you but I have to go to my next class. Try to @ a qualified helper :(

OpenStudy (anonymous):

so would the x^2 be a, the -7x be b and the 5 be c?

OpenStudy (anonymous):

oh okay thanks for your help

OpenStudy (anonymous):